# 0Year 0Year 0Year

### 10 Mar 2026 v2

@@ Abstract
Motivation: Phenotype concept recognition (CR) is a fundamental task in biomedical text mining. However, existing methods either require ontology-specific training making them struggle to generalize across diverse text styles and evolving biomedical terminology, or depend on general-purpose large language models (LLMs) that lack necessary domain knowledge.
@@ 1 Introduction
Biomedical text mining plays a key role in unlocking clinical and scientific knowledge from unstructured data sources such as clinical notes and research articles. A fundamental step in this process is ontology-based concept recognition (CR), which aims to identify textual mentions of concepts defined in a given ontology from input text in a zero-shot setting. An ontology is a formal, structured representation of domain-specific knowledge curated by experts, consisting of standardized concepts with associated names, definitions, synonyms, and hierarchical relationships. Phenotype CR, a specific instance of ontology-based CR using the Human Phenotype Ontology (HPO) (Talapova et al., 2023), has become a central research focus. This is largely due to the availability of richly annotated datasets (Weissenbacher et al., 2023; Lobo et al., 2017; Anazi et al., 2017), yet the task remains challenging because of the specialized and rapidly evolving nature of biomedical terminology and ontologies. Phenotype CR plays a critical role in downstream biomedical applications. For example, genetic disease diagnostics require accurate identification of phenotype concepts in clinical notes (Labbé et al., 2023), and biomedical knowledge graph construction relies on robust CR from scientific literature to support integrative data analysis and knowledge discovery (Huang et al., 2024).
@@
Traditionally, CR can be divided into two stages: (1) entity extraction, which identifies text spans to be linked to ontology concepts; and (2) entity linking, which maps these spans to semantically corresponding ontology concepts. Early CR methods are primarily dictionary-based, which rely on lookup tables and string-matching techniques to identify concept mentions. While these approaches offer high precision, they suffer from low recall due to limited vocabulary coverage and inability to handle linguistic variations (Jonquet et al., 2009; Taboada et al., 2014). In recent years, researchers have increasingly turned to neural methods that fine-tune domain-specific pre-trained language models (PLMs), such as BioBERT (Lee et al., 2020), on the rich synonym sets provided by the ontology. By distilling knowledge from synonymous relationships, these models effectively mitigate the challenges of lexical heterogeneity. However, as they are trained against a fixed ontology, re-training is required to recognize new concepts, limiting their usability for frequently updated ontologies like HPO as well as generalizability to unseen ontologies. In contrast, large language models (LLMs), such as GPT-4 (Achiam et al., 2023), have demonstrated strong zero-shot learning capabilities, offering new possibilities for phenotype CR. Recent studies have shown that LLMs can effectively extract clinical information, where phenotype concepts abound, without domain-specific fine-tuning (Meoni et al., 2023). Despite their potential, LLMs face challenges related to factual consistency and reliability in knowledge-intensive tasks (Reese et al., 2024). To address these challenges, retrieval-augmented generation (Lewis et al., 2020, RAG) has emerged as an effective technique to improve LLM’s CR performance by incorporating relevant information retrieved from the ontology through semantic similarity-based retrieval mechanisms. However, existing RAG-based methods often rely on general-purpose components that fail to capture the nuanced semantics of highly specialized and dynamic biological concepts.
@@
In this study, we propose AutoPCR, an automated phenotype CR framework based on prompting, which consists of three sequential steps. First, to ensure that extracted entities are biologically meaningful with high coverage, AutoPCR employs a unified entity extraction approach including BioNER and syntax-based strategies. In the second step, AutoPCR retrieves candidate concepts using a domain-specific semantic similarity model. Although phenotype CR has been extensively studied, prior work has largely overlooked advances in distantly-supervised biomedical entity linking. AutoPCR adopts one of these models named SapBERT (Liu et al., 2021) to accurately identify semantically relevant concepts. In the final step, AutoPCR performs entity linking by prompting an LLM. Each entity and its associated candidate set are encoded into a structured prompt that includes the synonyms and definition of each candidate concept. This approach enables accurate disambiguation even without ontology-specific fine-tuning. To further enhance its capability in distinguishing similar concepts, we introduce a self-supervised fine-tuning strategy that enables AutoPCR to learn from difficult positive and negative examples drawn by SapBERT. Experiments demonstrate the following advantages of AutoPCR: (1) Superior and robust performance. AutoPCR achieves the best average and most robust performance across various benchmarks, changing the status qua that prompt-based methods are lagging behind dictionary-based and neural methods. (2) Inductive capability. AutoPCR maintains superior performance even without prior exposure to HPO, making it adaptable to the frequent updates of HPO. (3) Generalizability. AutoPCR transfers well to a new ontology without reconfiguration and supports rapid deployment within minutes, offering great potential for broad application.
@@ 2 Related work
Dictionary-based methods use lookup tables or string similarity measures to identify entities by exhaustively matching input spans against ontology concepts. Tools such as NCBO (Jonquet et al., 2009), OBO (Taboada et al., 2014), Doc2HPO (Liu et al., 2019), and ClinPhen (Deisseroth et al., 2019) exemplify this strategy. Recently, FastHPOCR (Groza et al., 2024b) utilizes groups of morphologically equivalent words generated by GPT-4 to address lexical variations.
Neural methods leverage deep learning architectures, such as convolutional neural networks (CNNs) and BERT (Devlin et al., 2019), with supervised learning on ontology-derived synonym pairs to improve performance. There are two common entity extraction strategies adopted by neural methods, serving as the foundation for subsequent entity linking. One line of methods, e.g., PhenoTagger (Luo et al., 2021) and PhenoTagger++ (Qi et al., 2024), adopts a syntax-based strategy that exhaustively generates entity spans after removing punctuation and function words. PhenoTagger predicts concepts from extracted entities by fine-tuning a BioBERT. PhenoTagger++ improves PhenoTagger by augmenting training data with LLM-generated synonyms and character edition, as well as introducing graph-based concept embeddings from the ontology structure. The other line of methods use off-the-shelf BioNER tools such as Stanza (Zhang et al., 2021) to identify biomedical segments and then exhaustively extract entities from them. NCR (Arbabi et al., 2019) encodes entities using pre-trained text embeddings followed by a CNN and matches them with learnable embeddings of ontology concepts after hierarchical aggregation. PhenoBERT (Feng et al., 2022) builds upon NCR with multiple CNNs for phenotype subcategories to retrieve candidate concepts and a trainable BioBERT to score each entity-concept pair. Recently, PhenoBCBERT and PhenoGPT (Yang et al., 2024) train Bio+ClinicalBERT (Alsentzer et al., 2019) and GPT-3 (Brown et al., 2020) on manually labeled data to improve rare disease recognition.
LLMs, such as GPT-4, enable phenotype CR without the need of ontology-specific training, allowing for rapid adaptation to evolving biomedical vocabularies. Prompt-based methods, such as Labbé et al. (2023) and Groza et al. (2024a), employs prompt engineering to directly extract biomedical concepts from text. REAL (Shlyk et al., 2024) in a RAG style relies on LLMs to extract biomedical entities and perform entity linking. Candidate concepts are retrieved for each entity using a general-purpose embedding model based on names and definitions, and then passed to an LLM for final linking.
@@
Distant supervision has emerged as a practical solution for biomedical entity linking, where manual annotation is costly and often infeasible. Given the Unified Medical Language System (Bodenreider, 2004, UMLS), a comprehensive biomedical ontology with over 4M concepts covering HPO, these methods leverage ontology-derived supervision, enabling large-scale training without labeled corpora. BioSyn (Sung et al., 2020) aligns entities with concept synonyms using a combination of character-level features and learnable embeddings. SapBERT fine-tunes a PubMedBERT (Gu et al., 2021) through contrastive learning on hard concept name-synonym pairs. KrissBERT (Zhang et al., 2022) incorporates concept context from PubMed Canese and Weis (2013) into contrastive learning and employs a cross-attention encoder to re-rank candidate concepts. Its inference strategy retains multiple context embeddings per concept, enabling context-aware matching. Recent work (Sasse et al., 2024) shows that synthetic entities generated by LLMs can further enhance linking performance, especially under distribution shifts, highlighting the complementary role of generative models in distant supervision.
@@ 3 Method
Figure 1: Architecture of AutoPCR. It performs concept recognition in three stages: entity extraction using BioNER and syntactic entity extraction (e.g., extracted entities $e_{1}$, $e_{2}$, and $e_{3}$), candidate concept retrieval via SapBERT initialized from PubMedBERT (e.g., retrieved concepts $c_{1}$ and $c_{2}$ for $e_{3}$), and entity linking through prompting an LLM (e.g., linked concept $c_{1}$).
@@
We employ a unified entity extraction approach that combines BioNER and syntax-based strategies, making it applicable to both free-form clinical notes and standardized biomedical abstracts. For BioNER, we follow PhenoBERT and process each sentence using Stanza’s BioNER model to extract biomedically relevant segments. To improve coverage, additional segments are generated by splitting on punctuation and conjunctions, and all resulting segments are used as windows for $n$-gram entity extraction ($n{\in}[2,10]$). To better cover free-form text, we complement BioNER with a syntactic entity extraction component, which identifies structurally coherent phrases through noun phrase extraction, coordinated phrase decomposition, and abbreviation recovery. For noun phrase extraction, we employ the Berkeley Neural Parser (Kitaev and Klein, 2018, benepar) to obtain constituent trees for each sentence and recursively extract noun phrases. To handle coordinated expressions such as “broad and high nasal bridge”, which corresponds to both “broad nasal bridge” and “high nasal bridge” concepts, we introduce a coordinated phrase decomposition mechanism. Using spaCy’s dependency parser (Honnibal et al., 2020), we identify conjuncts for each coordinated token and then determine its longest constituent not overlapping with its conjuncts. Higher-level noun phrases covering multiple such constituents serve as templates to produce decomposed entities, where only one coordinated token’s constituent is used at once. This approach, compared with PhenoBERT which only handles A+B coordination, is capable of generalizing to more complex coordinated structures. Finally, we integrate the abbreviation detector from scispaCy (Neumann et al., 2019) to identify abbreviations and replace them by corresponding long forms, in order to reduce mismatch between abbreviations and unrelated identical concepts.
@@
To facilitate the LLM-based entity linker to distinguish between highly similar concepts within the ontology, we introduce a self-supervised fine-tuning strategy. We first generate challenging training examples by leveraging the internal structure of the ontology and the retrieval capabilities of SapBERT. For each ontology concept, we perform synonym cleaning by lemmatizing all associated synonyms and merging those that become identical. For each unique lemmatized synonym $s$ belonging to a ground-truth concept $c_{gt}$, we use SapBERT to retrieve the top-$r$ non-repetitive candidate concepts. We then construct positive and negative training examples as follows: To create a positive candidate set $C_{\text{cand}}^{+}$, we check if $c_{gt}$ is present in the retrieved list. If it is, the list is preserved. If not, we replace the last concept in the list with $c_{gt}$ to ensure the ground truth is present among difficult distractors. Note that $s$ should be removed from the synonym list of $c_{gt}$ to avoid leakage. To create a negative candidate set $C_{\text{cand}}^{-}$, we ensure $c_{gt}$ is absent. If $c_{gt}$ is in the list, we remove it and append the (r+1)(r{+}1)-th retrieved concept to the end. If not, we retain the list as is. This procedure yields two candidate sets of length $r$ for each synonym, representing the most difficult cases for SapBERT to resolve. We transform these examples into training instances by applying the prompt template above with two modifications: (1) we remove the “confidence” and “UMLS synonyms” requirements from the system prompt to simplify the learning objective, and (2) we set the desired output to the ground-truth concept ID for positive examples and “None” for negative examples. By iterating through all concept synonyms, we generate a total of 42,110 training instances. We fine-tune the entity linker using QLoRA (Dettmers et al., 2023) for one epoch with $r=3$ candidates to achieve efficient adaptation, and name the resulting model $\text{AutoPCR}_{\text{FT}}$ from now on.
@@ 4 Experiments
4.1 Benchmarks and baselines
To evaluate the performance of AutoPCR, we conduct experiments on four widely-used datasets: BIOC-GS, the development set from BioCreative VIII Track 3 (Weissenbacher et al., 2023), which consists of 382 clinical observation records from dysmorphology physical examinations, with an average length of 8.5 words. An example record is “ABDOMEN: Small umbilical hernia. Mild distention. Soft.” GSC-2024, a refined GSC+ (Lobo et al., 2017) dataset given by FastHPOCR, which comprises 228 PubMed abstracts, with an average length of 150 words. We use 22 of them for development, following PhenoTagger’s data split. ID-68, 68 real clinical notes from families with intellectual disabilities (Anazi et al., 2017), with an average length of 157 words. It is manually annotated by PhenoBERT in the same way as GSC+. NCBI (Doğan et al., 2014), consisting of 100 PubMed abstracts, with an average length of 205 words. BIOC-GS, GSC-2024, and ID-68 are grounded on HPO release 2024-02-08 under the root node “phenotypic abnormality” with 17,542 concepts. NCBI is grounded on MEDIC (Davis et al., 2025), a curated disease ontology integrating MeSH (Lipscomb, 2000) and OMIM (Amberger et al., 2019), under the root node “diseases” with 13,315 concepts. These datasets involve both free-form and standardized concept mentions, covering a wide range of use cases.
@@ 4.1 Benchmarks and baselines
We compare AutoPCR with competitive baselines from three categories: (1) dictionary-based methods, including NCBO, OBO, ClinPhen, and FastHPOCR; (2) neural methods, including NCR, PhenoTagger, PhenoBERT, and PhenoTagger++; and (3) prompt-based methods, including vanilla prompting (Groza et al., 2024a) and REAL. All baselines are tested with their recommended parameters. Prompt-based methods employ GPT-4o-mini as the LLM backend.
4.2 Implementation details
For entity extraction, we use Stanza packages “ner-i2b2” and “mimic”, spaCy model “en_core_web_trf”, benepar model “benepar_en3_large”, and scispaCy model “en_core_sci_scibert”. We recover abbreviations on ID-68, as they are counted separately. For candidate concept retrieval, we set $\tau_{1}{=}0.95{\in}\{0.9,0.925,0.95,0.975,1\}$ and $\tau_{2}{=}0.85{\in}\{0.8,0.825,0.85,0.875,0.9\}$ to distinguish high- and low-confidence matches. For similarity scores in $\tau_{2},\tau_{1})$, we retain up to $k{=}5{\in}\{3,5,7,9\}$ candidates. These hyperparameters are tuned on the GSC-2024 development set. For entity linking, we use Qwen3-Next-80B-A3B-Instruct as the frozen LLM backend and Qwen3-30B-A3B-Instruct-2507 as the base for fine-tuning (Yang et al., 2025). Temperature is fixed at zero for stability.
4.3 Main results
@@
Table 2: Document-level results on three datasets. AutoPCR is the highest in average F1 and F1 ranking.
Method BIOC-GS GSC-2024 ID-68 Avg. F1 Avg. Rk.
P R F1 (Rk.) P R F1 (Rk.) P R F1 (Rk.)
Dictionary- Based NCBO 82.64 45.70 58.85 (8) 99.33 52.26 68.49 (7) 88.75 64.69 74.84 (7) 67.39 7.33
OBO 73.93 42.72 54.15 (10) 85.98 52.40 65.12 (8) 82.84 60.28 69.78 (9) 63.01 9.00
ClinPhen 64.78 51.16 57.17 (9) 86.20 44.99 59.12 (10) 74.96 61.92 67.82 (10) 61.37 9.67
FastHPOCR 70.92 59.77 64.87 (5) 95.15 77.61 85.49 (1) 87.75 71.38 78.72 (3) 76.36 3.00
Neural NCR 65.54 57.62 61.32 (7) 81.88 73.09 77.24 (6) 79.18 77.68 78.42 (4) 72.33 5.67
PhenoTagger 71.33 65.89 68.50 (2) 87.57 76.62 81.73 (4) 84.16 72.38 77.83 (5) 76.02 3.67
PhenoBERT 76.53 53.97 63.30 (6) 90.33 73.87 81.27 (5) 94.11 78.56 85.64 (1) 76.74 4.00
PhenoTagger++ 69.53 64.24 66.78 (4) 89.38 76.70 82.55 (3) 79.81 72.76 76.12 (6) 75.15 4.33
Prompt- Based Vanilla 4.42 3.48 3.89 (11) 18.25 9.75 12.71 (11) 20.04 11.98 15.00 (11) 10.53 11.00
REAL 76.47 60.27 67.41 (3) 80.21 54.66 65.02 (9) 76.64 66.20 71.04 (8) 67.82 6.67
AutoPCR 84.57 67.44 75.04 (1) 89.36 78.32 83.48 (2) 84.33 76.71 80.34 (2) 79.62 1.67
$\text{AutoPCR}_{\text{FT}}$ 83.94 65.18 73.38 91.62 76.69 83.50 88.77 77.05 82.49 79.79
@@
The mention- and document-level results across three datasets grounded on HPO are shown in Tables 1 and 2, respectively. The two evaluation levels yield identical method rankings. AutoPCR ranks first on BIOC-GS and second on both GSC-2024 and ID-68, achieving the highest average F1 and the best overall rank. It also exhibits the most stable performance, demonstrating strong robustness across datasets. FastHPOCR achieves the strongest performance among dictionary-based methods, with the second-best average rank. Nevertheless, its performance drops notably on BIOC-GS, indicating reduced robustness on noisier clinical narratives. Traditional alternatives such as NCBO, OBO, and ClinPhen maintain high precision but suffer from low recall, resulting in overall lower F1. Neural methods PhenoTagger and PhenoBERT rank third and fourth overall. PhenoTagger performs better on BIOC-GS but worse on ID-68, whereas PhenoBERT shows the opposite trend. Both methods exhibit higher variability compared with AutoPCR. The earlier NCR is outperformed by its successor PhenoBERT, as expected, while PhenoTagger++, intended as a refinement of PhenoTagger, shows only a slight improvement on GSC-2024 but degraded performance on BIOC-GS and ID-68, leading to an overall decline and suggesting possible overfitting. Vanilla prompting performs poorly compared with all other baselines. In contrast, the RAG-based REAL performs strongly on BIOC-GS but ranks near the bottom on the others, highlighting the limitations of its generalist model design when applied to long, standardized text. These results comprehensively answer (RQ-1), demonstrating the superior and robust performance of AutoPCR across datasets. We further analyze these outcomes in detail, explaining how the design of AutoPCR contributes to its performance under different dataset characteristics.

### 02 Apr 2026 v3

@@ Abstract
Motivation: Phenotype concept recognition (CR) is a fundamental task in biomedical text mining. However, existing methods either require ontology-specific training, making them struggle to generalize across diverse text styles and evolving biomedical terminology, or depend on general-purpose large language models (LLMs) that lack necessary domain knowledge.
@@
Availability and Implementation: Our code is available at https://github.com/yctao7/AutoPCR.
Contact: drjieliu@umich.edu
@@ 1 Introduction
Biomedical text mining plays a key role in unlocking clinical and scientific knowledge from unstructured data sources such as clinical notes and research articles. A fundamental step in this process is ontology-based concept recognition (CR), which aims to identify textual mentions of concepts defined in a given ontology from input text in a zero-shot setting. An ontology is a formal, structured representation of domain-specific knowledge curated by experts, consisting of standardized concepts with associated names, definitions, synonyms, and hierarchical relationships. Phenotype CR, a specific instance of ontology-based CR using the Human Phenotype Ontology (HPO) (gargano2024mode), has become a central research focus. This is largely due to the availability of richly annotated datasets (weissenbacher2023phenoid; lobo2017identifying; anazi2017expanding), yet the task remains challenging because of the specialized and rapidly evolving nature of biomedical terminology and ontologies. Phenotype CR plays a critical role in downstream biomedical applications. For example, genetic disease diagnostics require accurate identification of phenotype concepts in clinical notes (labbe2023chatgpt), and biomedical knowledge graph construction relies on robust CR from scientific literature to support integrative data analysis and knowledge discovery (huang2024building).
@@
Traditionally, CR can be divided into two stages: (1) entity extraction, which identifies text spans to be linked to ontology concepts; and (2) entity linking, which maps these spans to semantically corresponding ontology concepts. Early CR methods are primarily dictionary-based, which rely on lookup tables and string-matching techniques to identify concept mentions. While these approaches offer high precision, they suffer from low recall due to limited vocabulary coverage and inability to handle linguistic variations (jonquet2009ncbo; taboada2014automated). In recent years, researchers have increasingly turned to neural methods that fine-tune domain-specific pre-trained language models (PLMs), such as BioBERT (lee2020biobert), on the rich synonym sets provided by the ontology. By distilling knowledge from synonymous relationships, these models effectively mitigate the challenges of lexical heterogeneity. However, as they are trained against a fixed ontology, re-training is required to recognize new concepts, limiting their usability for frequently updated ontologies like HPO as well as generalizability to unseen ontologies. In contrast, large language models (LLMs), such as GPT-4 (achiam2023gpt), have demonstrated strong zero-shot learning capabilities, offering new possibilities for phenotype CR. Recent studies have shown that LLMs can effectively extract clinical information, where phenotype concepts abound, without domain-specific fine-tuning (meoni2023large). Despite their potential, LLMs acquire ontology knowledge imperfectly during pretraining and cannot reliably disambiguate among the large number of highly specialized concepts in HPO without explicit ontology access (labbe2023chatgpt; groza2024evaluation). To address these challenges, retrieval-augmented generation (lewis2020retrieval, RAG) has emerged as an effective technique to improve LLM’s CR performance by incorporating relevant information retrieved from the ontology through semantic similarity-based retrieval mechanisms. However, the existing RAG-based method (shlyk2024real) relies on general-purpose components that fail to capture the nuanced semantics of highly specialized and dynamic biological concepts.
@@
In this study, we propose AutoPCR, an automated phenotype CR framework based on prompting, which consists of three sequential steps. First, AutoPCR employs a unified entity extraction approach including BioNER and syntax-based strategies with coordinated phrase decomposition, to ensure that extracted entities are biologically meaningful with broad coverage across both free-form clinical notes and standardized biomedical abstracts. In the second step, AutoPCR retrieves candidate concepts using a domain-specific semantic similarity model. Although phenotype CR has been extensively studied, prior work has largely overlooked advances in distantly-supervised biomedical entity linking. AutoPCR adopts one of these models named SapBERT (liu-etal-2021-self) to accurately identify semantically relevant concepts. AutoPCR then applies confidence-threshold routing that directly accepts high-confidence matches while delegating ambiguous cases to further LLM-based linking, improving both precision and recall. In the final step, AutoPCR performs entity linking by prompting an LLM. Each entity and its associated candidate set are encoded into a structured prompt that includes the synonyms and definition of each candidate concept. This approach enables accurate disambiguation even without ontology-specific fine-tuning. To further enhance its capability in distinguishing similar concepts, we introduce an optional self-supervised fine-tuning strategy that automatically constructs hard positive and negative training examples from ontology structure via SapBERT retrieval, enabling efficient ontology-specific adaptation without manual annotation. Together, AutoPCR introduces three technical novelties: a unified entity extraction approach that combines BioNER and syntax-based strategies, a domain-specific retrieval pipeline with confidence-threshold routing, and a self-supervised strategy for annotation-free ontology adaptation.
Experiments demonstrate the following advantages of AutoPCR: (1) Best average and robust performance across various benchmarks, changing the status quo that prompt-based methods are lagging behind dictionary-based and neural methods. (2) Inductive capability. AutoPCR maintains strong performance even without prior exposure to HPO, making it adaptable to the frequent updates of HPO. (3) Generalizability. AutoPCR transfers well to a new ontology without reconfiguration and supports rapid deployment within minutes, offering great potential for broad application.
@@ 2 Related work
Dictionary-based methods use lookup tables or string similarity measures to identify entities by exhaustively matching input spans against ontology concepts. Tools such as NCBO (jonquet2009ncbo), OBO (taboada2014automated), Doc2HPO (liu2019doc2hpo), and ClinPhen (deisseroth2019clinphen) exemplify this strategy. Recently, FastHPOCR (groza2024fasthpocr) utilizes groups of morphologically equivalent words generated by GPT-4 to address lexical variations.
Neural methods leverage deep learning architectures, such as convolutional neural networks (CNNs) and BERT (devlin2019bert), with supervised learning on ontology-derived synonym pairs to improve performance. There are two common entity extraction strategies adopted by neural methods, serving as the foundation for subsequent entity linking. One category of methods, e.g., PhenoTagger (luo2021phenotagger) and PhenoTagger++ (qi2024improved), adopts a syntax-based strategy that exhaustively generates entity spans after removing punctuation and function words. PhenoTagger predicts concepts from extracted entities by fine-tuning a BioBERT. PhenoTagger++ improves PhenoTagger by augmenting training data with LLM-generated synonyms and character edition, as well as introducing graph-based concept embeddings from the ontology structure. The other category of methods uses off-the-shelf BioNER tools such as Stanza (zhang2021biomedical) to identify biomedical segments and then exhaustively extract entities from them. NCR (arbabi2019identifying) encodes entities using pre-trained text embeddings followed by a CNN and matches them with learnable embeddings of ontology concepts after hierarchical aggregation. PhenoBERT (feng2022phenobert) builds upon NCR with multiple CNNs for phenotype subcategories to retrieve candidate concepts and a trainable BioBERT to score each entity-concept pair. Recently, PhenoBCBERT and PhenoGPT (yang2024enhancing) train Bio+ClinicalBERT (alsentzer2019publicly) and GPT-3 (brown2020language) on manually labeled data to improve rare disease recognition.
LLMs, such as GPT-4, enable phenotype CR without the need of ontology-specific training, allowing for rapid adaptation to evolving biomedical vocabularies. Prompt-based methods, such as labbe2023chatgpt and groza2024evaluation, employ prompt engineering to directly extract biomedical concepts from text. REAL (shlyk2024real) in a RAG style relies on LLMs to extract biomedical entities and perform entity linking. Candidate concepts are retrieved for each entity using a general-purpose embedding model based on names and definitions, and then passed to an LLM for final linking.
@@
Distant supervision has emerged as a practical solution for biomedical entity linking, where manual annotation is costly and often infeasible. Given the Unified Medical Language System (bodenreider2004unified, UMLS), a comprehensive biomedical ontology with over 4M concepts covering HPO, these methods leverage ontology-derived supervision, enabling large-scale training without labeled corpora. BioSyn (sung-etal-2020-biomedical) aligns entities with concept synonyms using a combination of character-level features and learnable embeddings. SapBERT fine-tunes a PubMedBERT (gu2021domain) through contrastive learning on hard concept name-synonym pairs. KrissBERT (zhang-etal-2022-knowledge) incorporates concept context from PubMed canese2013pubmed into contrastive learning and employs a cross-attention encoder to re-rank candidate concepts. Its inference strategy retains multiple context embeddings per concept, enabling context-aware matching. Recent work (sasse2024disease) shows that synthetic entities generated by LLMs can further enhance linking performance, especially under distribution shifts, highlighting the complementary role of generative models in distant supervision.
@@ 3 Method
Figure 1: Architecture of AutoPCR. It performs concept recognition in three stages: entity extraction using BioNER and syntactic entity extraction (e.g., extracted entities $e_{1}$, $e_{2}$, and $e_{3}$), candidate concept retrieval via SapBERT initialized from PubMedBERT (e.g., retrieved concepts $c_{1}$, $c_{2}$, $c_{3}$, and $c_{4}$ for $e_{3}$), and entity linking through prompting an LLM (e.g., linked concept $c_{2}$).
@@
We employ a unified entity extraction approach that combines BioNER and syntax-based strategies, making it applicable to both free-form clinical notes and standardized biomedical abstracts. For BioNER, we follow PhenoBERT and process each sentence using Stanza’s BioNER model to extract biomedically relevant segments. To improve coverage, additional segments are generated by splitting on punctuation and conjunctions, and all resulting segments are used as windows for $n$-gram entity extraction ($n{\in}[2,10]$). To better cover free-form text, we complement BioNER with a syntactic entity extraction component, which identifies structurally coherent phrases through noun phrase extraction, coordinated phrase decomposition, and abbreviation recovery. For noun phrase extraction, we employ the Berkeley Neural Parser (kitaev2018constituency, benepar) to obtain constituent trees for each sentence and recursively extract noun phrases. To handle coordinated expressions such as “broad and high nasal bridge”, which corresponds to both “broad nasal bridge” and “high nasal bridge” concepts, we introduce a coordinated phrase decomposition mechanism. Using spaCy’s dependency parser (honnibal2020spacy), we identify conjuncts for each coordinated token and then determine its longest constituent not overlapping with its conjuncts. Higher-level noun phrases covering multiple such constituents serve as templates to produce decomposed entities, where only one coordinated token’s constituent is used at once. This approach, compared with PhenoBERT which only handles A+B coordination, is capable of generalizing to more complex coordinated structures. Finally, we integrate the abbreviation detector from scispaCy (neumann-etal-2019-scispacy) to identify abbreviations and replace them by corresponding long forms, in order to reduce mismatch between abbreviations and unrelated identical concepts.
@@
To facilitate the LLM-based entity linker to distinguish between highly similar concepts within the ontology, we introduce a self-supervised fine-tuning strategy. We first generate challenging training examples by leveraging the internal structure of the ontology and the retrieval capabilities of SapBERT. For each ontology concept, we perform synonym cleaning by lemmatizing all associated synonyms and merging those that become identical. For each unique lemmatized synonym $s$ belonging to a ground-truth concept $c_{gt}$, we use SapBERT to retrieve the top-$r$ non-repetitive candidate concepts. We then construct positive and negative training examples as follows: To create a positive candidate set $C_{\text{cand}}^{+}$, we check if $c_{gt}$ is present in the retrieved list. If it is, the list is preserved. If not, we replace the last concept in the list with $c_{gt}$ to ensure the ground truth is present among difficult distractors. Note that $s$ should be removed from the synonym list of $c_{gt}$ to avoid leakage. To create a negative candidate set $C_{\text{cand}}^{-}$, we ensure $c_{gt}$ is absent. If $c_{gt}$ is in the list, we remove it and append the (r+1)(r{+}1)-th retrieved concept to the end. If not, we retain the list as is. This procedure yields two candidate sets of length $r$ for each synonym, representing the most difficult cases for SapBERT to resolve. We transform these examples into training instances by applying the prompt template above with two modifications: (1) we remove the “confidence” and “UMLS synonyms” requirements from the system prompt to simplify the learning objective, and (2) we set the desired output to the ground-truth concept ID for positive examples and “None” for negative examples. By iterating through all concept synonyms, we generate a total of 42,110 training instances. We fine-tune the entity linker using QLoRA (dettmers2023qlora) for one epoch with $r=3$ candidates to achieve efficient adaptation, and name the resulting model $\text{AutoPCR}_{\text{FT}}$ from now on.
@@ 4 Experiments
Benchmarks and baselines
To evaluate the performance of AutoPCR, we conduct experiments on four widely-used datasets: BIOC-GS, the development set from BioCreative VIII Track 3 (weissenbacher2023phenoid), which consists of 382 clinical observation records from dysmorphology physical examinations (avg. 8.5 words, e.g., “ABDOMEN: Small umbilical hernia. Mild distention. Soft.”). GSC-2024, a refined GSC+ (lobo2017identifying) dataset given by FastHPOCR, which comprises 228 PubMed abstracts (avg. 150 words). We use 22 of them for development and the rest for evaluation, following PhenoTagger’s data split. ID-68, 68 real clinical notes (anazi2017expanding) from families with intellectual disabilities (avg. 157 words). It is manually annotated by PhenoBERT in the same way as GSC+. No splits are provided so we use all of them for evaluation. NCBI (dougan2014ncbi), of which we use the test split of 100 PubMed abstracts (avg. 205 words). The ontologies used are HPO release 2024-02-08 under the root “phenotypic abnormality” (17,542 concepts) for BIOC-GS, GSC-2024, and ID-68, and MEDIC (davis2025comparative), a curated disease ontology integrating MeSH (lipscomb2000medical) and OMIM (amberger2019omim), under the root “diseases” (13,315 concepts) for NCBI. These datasets involve both free-form and standardized concept mentions, covering a wide range of use cases.
@@
We compare AutoPCR with competitive baselines from three categories: (1) dictionary-based methods, including NCBO, OBO, ClinPhen, and FastHPOCR; (2) neural methods, including NCR, PhenoTagger, PhenoBERT, and PhenoTagger++; and (3) prompt-based methods, including Vanilla, using the best-performing zero-shot template (“Prompt 4”) from groza2024evaluation, and REAL. All baselines are tested with their recommended parameters. Prompt-based methods employ GPT-4o-mini as the LLM backend.
Implementation details
For entity extraction, we use Stanza packages “ner-i2b2” and “mimic”, spaCy model “en_core_web_trf”, benepar model “benepar_en3_large”, and scispaCy model “en_core_sci_scibert”. We recover abbreviations on ID-68, as they are counted separately. For candidate concept retrieval, we set $\tau_{1}{=}0.95{\in}\{0.9,0.925,0.95,0.975,1\}$ and $\tau_{2}{=}0.85{\in}\{0.8,0.825,0.85,0.875,0.9\}$ to distinguish high- and low-confidence matches. For similarity scores in $\tau_{2},\tau_{1})$, we retain up to $k{=}5{\in}\{3,5,7,9\}$ candidates. These hyperparameters are tuned on the GSC-2024 development set. For entity linking, we use Qwen3-Next-80B-A3B-Instruct as the frozen LLM backend and Qwen3-30B-A3B-Instruct-2507 as the base for fine-tuning (yang2025qwen3). Temperature is fixed at zero for stability.
Main results
@@
The mention- and document-level results across three datasets grounded on HPO are shown in Tables 1 and Supplementary Table 1, and visualized in Supplementary Figures 1 and 2, respectively. The two evaluation levels yield identical method rankings. AutoPCR ranks first on BIOC-GS and second on both GSC-2024 and ID-68, achieving the highest average F1 and the best overall rank. It also exhibits the most stable performance, demonstrating strong robustness across datasets. FastHPOCR achieves the strongest performance among dictionary-based methods, with the second-best average rank. Nevertheless, its performance drops notably on BIOC-GS, indicating reduced robustness on noisier clinical narratives. Traditional alternatives such as NCBO, OBO, and ClinPhen maintain high precision but suffer from low recall, resulting in overall lower F1. Neural methods PhenoTagger and PhenoBERT rank third and fourth overall. PhenoTagger performs better on BIOC-GS but worse on ID-68, whereas PhenoBERT shows the opposite trend. Both methods exhibit higher variability compared with AutoPCR. The earlier NCR is outperformed by its successor PhenoBERT, as expected, while PhenoTagger++, intended as a refinement of PhenoTagger, shows only a slight improvement on GSC-2024 but degraded performance on BIOC-GS and ID-68, leading to an overall decline and suggesting possible overfitting. Vanilla prompting performs poorly compared with all other baselines. In contrast, the RAG-based REAL performs strongly on BIOC-GS but ranks near the bottom on the others, highlighting the limitations of its generalist model design when applied to long, standardized text. These results comprehensively answer (RQ-1), demonstrating the strong and robust performance of AutoPCR across datasets. We further analyze these outcomes in detail, explaining how the design of AutoPCR contributes to its performance under different dataset characteristics.